Weather Report for March 31, 2011

Cirrus slowed observations often during the first half of the night. After midnight, the cloud cover became too thick to continue. After two hours, the dome could be reopened. There was little photometry attempted tonight, but about a dozen blazars were successfully observed for polarization. The seeing was 1-2 arcsec during the useful portions of the night and was quite good late. Monitoring will start again at the beginning of scheduled campaign 27 on April 4, 2011.